OpenStudy (anonymous):

Find the inverse of the function y = 4x – 7. Is the inverse a function?

OpenStudy (anonymous):

yes. y=(x+7)/4

OpenStudy (anonymous):

inverse is: y = (x + 7) / 4 or y = 1/4x + 7/4

OpenStudy (anonymous):

same result. understand?